# RIGNO: A Graph-based framework for robust and accurate operator learning for PDEs on arbitrary domains

### v1

@@ F.1 CNO
Up/downsampling layers: $4$
Residual blocks in the bottleneck: $6$
Residual blocks in the middle layers: $6$
@@ F.4 Geo-FNO and FNO DSE
Number of Fourier layers: $4$
@@ F.5 GINO
Number of Fourier layers: $3$
@@
Appendix G Further experimental results
There are many hyperparameters that can affect the training stability as well as different aspects of the performance of a trained RIGNO. In the experiments presented in this work, we fix a set of architectural choices that perform fairly well for all the datasets and training sizes. These hyperparameters are listed in Table E.1. Note that these hyperparameters do not correspond to the best-performing architecture, but rather they guarantee a fair performance with all the desired properties. In Section G.2, we show that the performance can be improved by using larger models. In the experiments of this chapter, we perturb one or a few of these hyperparameters while leaving the rest fixed in order to study how the performance of the model changes with respect to the changed hyperparameters. The trainings are carried out using the settings described in Section E, unless otherwise stated.
G.1 Training Randomness
In order to quantify the dependence of the final results on training randomness, we train RIGNO five times, each time with a different seed for the pseudo-random number generator. We store the model with the lowest validation error during training as in all experiments, and check for fluctuations in the test error across all five repetitions of each experiment. In these experiments, we use the unstructured version of the datasets (random point clouds), we train using 512 training samples with batch size 8, and do not fine tune with fractional pairing. The results are summarized in Table G.1, where the maximum and minimum test errors across the repetitions are reported for each dataset. These fluctuations are attributed to the randomness in processes such as weight initializations, dataset shuffling, edge masking, and random sub-sampling of the coordinates. In the rest of the experiments, we only train each model once with a fixed training random seed, keeping in mind that the errors can slightly deviate because of training randomness.
Table G.1: Statistics of relative $L^{1}$ test errors with different random seeds. We train RIGNO-12 on 512 samples from each datasets. Each experiment is repeated 5 times.
Error [%]
Dataset Mean $\pm$ Standard deviation
NS-PwC 1.05 $\pm$ 0.040
NS-SVS 2.50 $\pm$ 0.157
G.2 Scaling with Model Size
In order to evaluate how the performance of RIGNO scales with model size, we start with RIGNO-12 as the base architecture and change the latent sizes (all at the same time), the number of message passing steps in the processor, the overlap factors (both at the same time), and the number of times that the regional mesh is sub-sampled when processor edges are built (number of regional edge levels). We train these models on 256 samples of the NS-PwC dataset without fractional pairing and calculate the error at $t_{14}$ on 128 test samples by using the AR-2 rollout scheme (defined in Section C.2). The results are plotted in Figure G.1.
We observe that the latent sizes and number of processor steps have the largest effect on the model capacity. The error consistently falls down when these hyperparameters are scaled, and we expect to get even lower errors with larger RIGNO architectures. With larger overlap factors, although the number of edges (and hence the computational cost) increases rapidly, the rate of convergence is relatively low. Therefore, the authors recommend using the minimum overlap, especially for the encoder. With more regional mesh levels, on the other hand, the number of edges only increases logarithmically, which is also reflected in the inference times. With 4 regional mesh levels, we obtain $3.75\%$ error, which is extremely lower than the $17.10\%$ error with only a single regional mesh level. The inference time, however, is only increased by $14\%$. These results show the effectiveness of longer edges in the construction of the regional mesh. However, depending on the space resolution of the inputs, there is saturation limit for this hyperparameter. If we consider a downsampling factor of 4, with $00016^{\prime}$ 000 input coordinates, for instance, we end up with a mesh with only $4$ nodes after 6 downsampling steps. With so few nodes we can only get 6 additional edges at most, which explains why the decreasing trend in the error stops at 4 downsampling steps.
Figure G.1: Test errors at $t=t_{14}$ with different strategies for scaling model size. The experiments are carried out on the NS-PwC dataset.
G.3 Temporal Continuity: Generalization to Unseen Lead Times
In order to assess the generalization of the trained models to unseen lead times $\tau$, we train RIGNO on the NS-PwC dataset with $2\Delta t$ time resolution, and test it on the test dataset with $\Delta t$ time resolution. With the incorporated mechanisms, namely time-conditioned normalization layers, all2all training, and fractional pairing fine-tuning, we expect to see a monotone increase in the error as we increase the lead time; which is very well captured in our tests as visualized in Figure G.2. Here, the lead time $\tau$ is fixed at each line in the plot, and the input time snapshot $t$ is changed. This explains why the first entries are missing for larger lead times. In the uppermost line ($\tau=10\Delta t$), for instance, the first entry corresponds to $t=0$, $\tau=10\Delta t$; the second entry to $t=1\Delta t$, $\tau=10\Delta t$; and so on.
In Figure G.2, the errors increase with larger lead times, which is expected since estimating the solution of the PDE with a higher lead time is inherently more difficult. In these experiments, we have removed the pairs with $\tau>8\Delta t$ from the all2all training pairs. Hence, $\tau=9\Delta t$ and $\tau=10\Delta t$ are considered as extrapolation in lead time. Since lead times $\tau=3\Delta t$, $\tau=5\Delta t$, and $\tau=7\Delta t$ are also not present in the all2all pairs, they are considered as interpolation in lead time. We can see that the monotonicity of the increase in the error is preserved for the extrapolated as well as the interpolated lead times. Moreover, the error increases in a controlled manner. The authors believe that this property is attributed to incorporation of time-conditioned normalization layers in the architecture and the use of time-derivative time-stepping strategy. However, with the smallest lead time $\tau=\Delta t$, it was observed that the monotonicity was not preserved without fine-tuning with the fractional pairing strategy.
Figure G.2: Direct test errors with different $\tau$ for the NS-PwC dataset. The vertical line indicates $t=t_{14}$, which is the last time snapshot seen by the model during training. Different lines denote different lead times. The gray lines correspond to lead times present in the training dataset; the green lines highlight interpolation in lead time; and the red lines indicate extrapolation in lead time. The model is trained on 512 samples with batch size 4.
G.4 Model Uncertainty
The edge masking technique, randomly disables a subset of the edges in each message-passing block. By adding randomness to the model, edge masking allows inferring the same model multiple times with the same input in order to generate an ensemble of estimates at a later time. The standard deviation of the generated ensemble can be interpreted as an indicator of model uncertainty on each output coordinate.
Figure G.3 visualizes the mean and the standard deviation of an ensemble of estimates at $t=t_{14}$ generated from input at $t=t_{0}$, along with the absolute error of the mean with respect to the ground-truth solution. More such visualizations are available in Section H. These visualizations show how the regions with the highest errors largely overlap with the regions where the ensemble has the highest deviations; which supports the interpretation of the deviations as a measure for model uncertainty, and shows that they are good indicators for inaccuracies in the estimates generated by the model.
Figure G.3: Mean and standard deviation of an ensemble of autoregressive estimates at $t_{14}$ for a test sample of the NS-SL dataset.
G.5 Robustness to Noise
(a) Heat-L-Sines
(b) NS-PwC
Figure G.4: Autoregressive test errors introduced by the noise in the input (relative). The y-axes show the difference in the relative error with noisy and clean inputs. The solid lines (best) show the results with edge masking and overlap factors 2.0. The dashed lines show the results with the same overlap factors but without edge masking during training. The dotted lines show the results with edge masking and overlap factors 1.0. All models are trained on 256 trajectories, without all2all training, and without fractional pairing fine-tuning.
For each input in the test dataset, we compute its standard deviation $\sigma_{\rm input}$, and add a Gaussian noise with standard deviation $\sigma_{\rm noise}=C_{noise}\sigma_{\rm input}$, with a noise level $C_{noise}\in\{0.005,0.02\}$. Starting with a noisy input, we apply the operators autoregressively with $\tau_{\rm max}=2\Delta t$. The errors in the resulting trajectories are partially due to the input noise, and partially to the approximation error of the operator. In order to distinguish between these two, we also unroll the operator with clean input ($C_{noise}=0$), and subtract the resulting errors from the errors with noisy inputs. This procedure gives a measure for the part of the error that is introduced by the noise in the input, denoted by noise-induced error. We present the relative noise induced error with respect to the approximation error: $|{e_{\rm noisy}-e_{\rm clean}}|/{e_{\rm clean}}$, where $e$ represents the autoregressive test error at $t=t_{14}$. The results are visualized in Figure G.4. With $2\%$ input noise at $t_{0}$, the noise is very well dampened to only $0.5\%$ at $t_{14}$ of the Heat-L-Sines dataset. With larger support sub-regions, the regional nodes aggregate data from a larger pool of physical nodes. Therefore, they are less sensitive to noisy inputs, and a stronger noise dampening is observed. Edge masking also plays an important role in noise dampening. When larger support sub-regions are used along with edge masking, $2\%$ input noise is dampened below $0.2\%$ at $t_{14}$.
@@ Appendix H Visualizations
Model estimates and their absolute errors, along with model input and the ground-truth output are visualized in Figures H.1-H.20. Visual inspection of model estimates for the CE-Gauss and CE-RP datasets reveals that there are two main reasons that explain the high errors for these datasets. First, model estimates are generally smoother than the ground-truth solutions. This inability to produce high-frequency estimates is reflected in the high errors where vortices are present. This is evident in the high-frequency details of velocity components in these model estimates. Second, although the model is able to capture shocks quite well at the right locations, it generates a smoother transition with smaller gradients. This is very well witnessed in the absolute estimate errors of Figure H.18, where the highest errors are located in small neighborhoods around shocks. The statistics of ensembles of estimates are visualized in Figures H.21-H.24. Estimate ensembles are generated by unrolling the operator randomly 20 times. Because of random edge masking, the estimates will slightly be different every time. The statistics of the estimate ensembles can act as indicators for model uncertainty. See Section G.4 for a broader discussion.

### v2

@@ G.1 CNO
Up/downsampling layers: 44
Residual blocks in the bottleneck: 66
Residual blocks in the middle layers: 66
@@ G.4 Geo-FNO and FNO DSE
Number of Fourier layers: 44
@@ G.5 GINO
Number of Fourier layers: 33
@@
G.6 MeshGraphNet
The MeshGraphNet [33] architecture is an end-to-end graph-based neural operator that performs message passing directly on the given coordinates (physical nodes). Since the architecture shares many similarities to RIGNO, we can obtain a MeshGraphNet by making a few modifications to our architecture. We consider RIGNO-12 and perform the downscaling (encoder) and upscaling (decoder) processes by considering very small overlap factors, i.e., very small support radii. Regional nodes lose their meaning since each of them is only connected to one corresponding physical node. Removing the multi-scale edges in this architecture leads to a MeshGraphNet with 1.9M parameters. We train this MeshGraphNet with edge masking to be more consistent with RIGNO but use a larger batch size (64) to compensate for the extra training time due to removing the regional nodes. All other hyperparameters and settings are the same as reported in Section E.
G.7 UPT
The Universal Physics Transformer (UPT) is a unified neural operator framework designed for flexible and scalable learning across a wide range of spatio-temporal problems, including both Eulerian (grid-based) and Lagrangian (particle-based) simulations [1]. UPT consists of three components: an encoder that compresses the input into a latent representation, a transformer-based approximator that propagates dynamics entirely in the latent space, and a decoder that queries this representation at arbitrary spatial locations. The model is conditioned on time and boundary conditions via DiT-style feature modulation.
We adopt a lightweight configuration of UPT in our experiments. Specifically, we use 64 latent tokens and 64 hidden dimensions. For time-dependent datasets, this results in 2.18M trainable parameters, and for time-independent datasets (without time conditioning), the number of parameters is 0.74M. A larger variant with 256 latent tokens and 192 hidden dimensions with 10.7M parameters has also been considered. We found that the large UPT strongly suffers from optimization difficulties on the datasets considered in this work, leading to higher errors than the ones reported in Table 1 with all datasets. In all the presented benchmarks, UPT architecture and model hyperparameters are fixed for all experiments as follows:
Trainable parameters: 2.182.18M or 0.740.74M (see above)
Number of epochs: $500$
Number of latent tokens: $64$
Hidden dimensions: $64$
Number of supernodes: $2048$
Encoder and decoder depth: 44 layers each
Approximator depth: 44 layers
Number of attention heads: 44
Optimizer: AdamW
Learning rate: $1\cdot 10^{-3}$
Scheduler: Cosine decay
Batch size: $32$
Early stopping: If validation loss does not improve for $40$ epochs
G.8 Transolver
Transolver [45] is an attention-based neural operator framework that learns global latent tokens from input features and processes them through self-attention mechanisms. In our experiments, we observed that this architecture scales poorly to large-scale datasets. This limitation arises from its use of global interactions implemented as independent MLPs between input points and sparse tokens. Consequently, a very large number of MLP evaluations are required, quickly exhausting GPU memory. For instance, on an RTX 4090 GPU, the maximum feasible Transolver model size during training is approximately 0.5M parameters. For time-independent datasets, we were able to increase the model size to 0.9M parameters. The corresponding results are summarized in Table G.1. While both RIGNO variants exhibit clear advantages on time-dependent problems, the accuracy is comparable for elliptic problems.
Table G.1: RIGNO and Transolver benchmarks on datasets with unstructured space discretizations. All training datasets have 1024 trajectories except for the AF dataset (2048).
Dataset Median relative $L^{1}$ error [%]
RIGNO-18 RIGNO-12 Transolver
Wave-C-Sines 5.35 6.25 27.7
CE-RP 3.98 4.92 17.8
Poisson-Gauss 2.26 2.52 2.02
AF 1.00 1.09 1.21
Elasticity 4.31 4.63 4.92
@@ Appendix H Visualizations
Model estimates and their absolute errors, along with model input and the ground-truth output are visualized in Figures H.1-H.21. Visual inspection of model estimates for the CE-Gauss and CE-RP datasets reveals that there are two main reasons that explain the high errors for these datasets. First, model estimates are generally smoother than the ground-truth solutions. This inability to produce high-frequency estimates is reflected in the high errors where vortices are present. This is evident in the high-frequency details of velocity components in these model estimates. Second, although the model is able to capture shocks quite well at the right locations, it generates a smoother transition with smaller gradients. This is very well witnessed in the absolute estimate errors of Figure H.19, where the highest errors are located in small neighborhoods around shocks. The statistics of ensembles of estimates are visualized in Figures H.22-H.25. Estimate ensembles are generated by unrolling the operator randomly 20 times. Because of random edge masking, the estimates will slightly be different every time. The statistics of the estimate ensembles can act as indicators for model uncertainty. See Section F.6 for a broader discussion.
@@
Figure H.9: Estimation of RIGNO-12 for a random test sample of the unstructured Poisson-Gauss dataset.